### v1

# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we first propose EventHallusion, a novel benchmark that focuses on assessing the VideoLMMs’ hallucination phenomenon on video event comprehension. Based on the observation that existing VideoLLMs are entangled with the priors stemming from their foundation models, our EventHallusion is curated by meticulously collecting videos and annotating questions to intentionally mislead the VideoLLMs into interpreting events based on these priors rather than accurately understanding the video content. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD suppresses the model’s preference toward their priors by comparing the original video with a constructed counterpart, whose temporal cues are disrupted, during the autoregressive decoding stage. Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we find that the open-source models suffer significantly from hallucination problems, whereas the closed-source models perform markedly better. By further equipping open-sourced VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023; Dettmers et al. 2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al. 2022; Liu et al. 2024a; Bai et al. 2023; Jiao et al. 2024b, a). Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (Jiao et al. 2024a; Yin et al. 2023b). Despite their great success, they still face hallucination problems, where models generate plausible yet factually incorrect responses (Leng et al. 2024), potentially threatening reliability and security in real-world applications. To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.  (Li et al. 2023b; Fu et al. 2023; Liu et al. 2023b).
However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  (Ma et al. 2023; Li, Wang, and Jia 2023). To mitigate this gap, two recent novel benchmarks (Wang et al. 2024; Guan et al. 2024) have been proposed. Specifically, as illustrated in Fig.1, HallusionBench (Guan et al. 2024) reverses all frames within a video and then queries the model to determine whether the event described in the original captions occurs in the altered video. VideoHallucer (Wang et al. 2024) employs long videos containing multiple events and then queries the model to assess the chronological order of these events. Generally, compared to those focused on hallucinations in static images, benchmarks that assess event hallucinations in dynamic videos lag behind.
When delving into VideoLLMs’ hallucinations on events, we found that the major causes can be attributed to language priors of LLMs (Agrawal, Batra, and Parikh 2016; Agarwal, Shetty, and Fritz 2020) and statistical bias of the inherited ImageLLM’s training corpus (Leng et al. 2024). As shown in Fig.1, when presented with a woman and a bike, Video-LLaVA (Lin et al. 2023), LLaVA-NeXT-Video (Li et al. 2024a), and ShareGPT4Video (Chen et al. 2024) associate them with the intuitive assumption of ”riding the bicycle”, rather than recognizing the clear fact of ”carrying the bicycle”. However, this essential event-related hallucination cannot be effectively assessed in both VideoHallucer (Wang et al. 2024) and HallusionBench (Guan et al. 2024). As shown in Fig.1, VideoHallucer has already provided the correct events in the question, merely requiring the model to determine their chronological orders without analyzing the events themselves. Although HallusionBench addresses event-oriented hallucinations, the frame reversal operations employed to create test samples may deviate from the natural distribution of videos.
Figure 1: Comparison of our proposed EventHallusion with previous benchmarks. Different from requiring models to locate events or comprehend reversed videos, our EventHallusion focuses on evaluating inherent priors-oriented hallucinations concerning events in VideoLLMs. Green and red represent the original and hallucination settings for evaluation in HallusionBench and VideoHallucer. At the bottom, we exhibit an example in our EventHallusion where Video-LLaVA, LLaVA-NeXT-Video, and ShareGPT4Video generate hallucinations towards the video event with hallucinated content marked by red.
Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates state-of-the-art VideoLLMs for hallucinations concerning events that are susceptible to the model’s inherent priors. The primary challenge in developing the EventHallusion is sourcing appropriate video data. Since existing publicly available video datasets, such as AcitivityNet (Caba Heilbron et al. 2015), randomly sample video fragments from daily life. As a result, their data distribution naturally reflects the priors resulting from the high-frequency co-occurrence of objects and event bindings (e.g., a person riding a bike in most cases). Therefore, instead of refactoring samples from existing video datasets, we manually collect videos and annotate question-answer pairs which may lead VideoLLMs to interpret events based on their pre-existing priors from prevalent internet video platforms. Specifically, to comprehensively probe existing VideoLLMMs’ sensitivity to their learned priors, we categorize our benchmark into two basic categories as shown in Fig.2: (1) In the first category, all videos are meticulously selected to ensure that the included events occur infrequently. Within this category, we further split videos into two groups according to their event composition. Videos in the first group (i.e., “entire rare events” in Fig.2) depict a rare event throughout the entire duration. And videos in the second group (i.e., “interleaved common-rare events” in Fig.2) include a period depicting a common event, followed by a rare event unexpectedly happens. Compared with those in the first group, videos in the second group are more challenging because the initially presented common event acts as a demonstration to consolidate the model’s tendency to their priors, which is also proved in Tab. 1. (2) In the second category, all questions are customized to include a plausible but incorrect common event as a distraction, and videos depict events that occur frequently. Instead of evaluating VideoLLM’s sensitivity to video content as in the first category, the testing samples in the second category are designed to assess the model’s sensitivity to question content, where the included misleading event descriptions can trigger the model to rely on its pre-existing priors. Through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination problems on our benchmark, which can provide valuable insights for further research.
Moreover, inspired by a hallucination elimination approach for image LLMs (i.e., VCD (Leng et al. 2024)), we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs. As previously analyzed, VideoLLMs can be easily influenced by their prior knowledge. Therefore, our TCD is designed to mitigate such biases and enhance the model’s objective judgments in event comprehension. Technically, for each given video, we first construct its counterpart with temporal downsampling and shuffling data augmentation operations. Both the original video and its counterpart are then fed into the VideoLLMs simultaneously to generate a pair of predicted logits, which are used for contrastive decoding similar to the approach in VCD (Leng et al. 2024). In the proposed TCD, the created counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s prior knowledge. Subtracting these amplified priors in the logits space during the contrastive decoding process helps guide the model toward more reliable judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We conduct comprehensive ablation studies to provide more in-depth analysis.
Figure 2: Illustration of evaluation categories. Our EventHallusion is basically divided into two categories: Sensitivity to Video Content and Sensitivity to Question Content. The content that induces VideoLLMs to generate event-related hallucinations is marked by red.
Figure 3: Framework of Temporal Contrastive Decoding. We construct the temporal modeling distorted video as the comparison group with fewer frames extracted and then contrast their logits to mitigate the event hallucination.
Related Works
Video Large Language Models. In the natural language processing field, Large Language Models (LLMs) have demonstrated astounding performance in solving traditional tasks. What’s more, LLMs also exhibit extraordinary generalization in handling self-defined tasks described by user’s instructions, making them capable of playing the role of versatile assistants. Benefiting from such abilities, Multimodal Language Models (MLLMs) have emerged by introducing visual inputs into LLMs, promoting the unified comprehension of both textual and visual content.
Video-supported MLLMs are further proposed since the established vision-text understanding ability of image MLLMs. To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward. Video-ChatGPT (Maaz et al. 2023) utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA (Zhang, Li, and Bing 2023) introduces a Video Q-Former to aggregate frame-level representations. Vista-LLaMA (Ma et al. 2023) preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality. LLaMA-VID (Li, Wang, and Jia 2023) proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information. However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
Video Evaluation Benchmarks. Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models. To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made (Fu et al. 2023; Liu et al. 2023b; Li et al. 2024c). Their evaluation comprises various aspects, including perception, reasoning, and cognition. With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities. Tempcompass (Liu et al. 2024b) focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order. FunQA (Xie et al. 2023) proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
Despite previous works sharing similarities with ours, there exist several critical differences. First, we focus on exploring and evaluating the hallucination phenomenon in the event aspects of existing VideoLLMs. Furthermore, we emphasize and validate that the event hallucination phenomenon of video LLMs is influenced by language prior. Second, based on this conclusion, we collect videos that are inclined to cause event hallucinations in VideoLLMs and construct the benchmark for evaluation. Third, we propose a Temporal Contrastive Decoding (TCD) to alleviate the event hallucination, which exhibits effectiveness in multiple VideoLLMs.
Hallucination in MLLMs. Hallucination is a phenomenon widely observed in MLLMs. The occurrence of hallucinations is attributed to various factors, including imbalanced training data distribution, weak attention to visual content, and other possible reasons. To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed. Woodpecker (Yin et al. 2023a) adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano (Lee et al. 2023) proposes a critique-revise-decide process to self-revise their initial generated answers. Besides, LRV-Instruction (Liu et al. 2023a) and HalluciDoctor (Yu et al. 2024) propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training. VCD (Leng et al. 2024) employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD (Zhu et al. 2024) constructs the comparison pair by introducing the image-biased and original MLLMs. Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs. Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
EventHallusion Benchmark
In this section, we first analyze the vulnerabilities of VideoLLMs emphasized by each category in our EventHallusion benchmark. Then, we elaborate on the detailed curation process of video data and question-answer pair annotation. Finally, we introduce the metrics calculation approach.
Vulnerability Dimension Analysis
To comprehensively assess a VideoLLM’s sensitivity to its priors across diverse aspects, we employ two categories to examine the influence of video and question content, respectively.
Sensitivity to Video Content
In this category, we intentionally collect videos depicting events that infrequently occur in daily life (e.g., a person carrying a bike, as shown in Fig.1). Such videos can easily trigger the prior knowledge learned from foundational ImageLLMs (noting that existing open-source VideoLLMs are all derived from certain ImageLLMs). As a result, the VideoLLM may rely on these strong visual priors rather than accurately interpreting the temporal cues of the video.
Sensitivity to Question Content
In this category, we select videos depicting events that commonly occur in daily life to exclude the influence of the visual priors mentioned previously. Additionally, we introduce a frequently occurring event that is inconsistent with the ground truth depicted in the video (e.g., “eating” in the Fig.2). Such customized questions can easily trigger priors of the LLM which obtains knowledge from enormous textual corpus. As a result, the VideoLLM may rely on its language priors to find a shortcut in answering the question.
Video Data Collection
When collecting video sources for our EventHallusion, we adhere to the following two rules to ensure high quality. Firstly, the objects and inter-object interaction details should be presented clearly to avoid visual detail missing. Secondly, the scene presented in the video is kept consistent without any shot cuts to avoid the adverse effects of environmental variation.
In total, we collect 400 eligible videos from multiple open video platforms on the Internet. To avoid data leakage, we have blurred the captions in the video, thereby encouraging models to analyze video events solely through visual content. These 400 videos cover 7 domains including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, and Nature as shown in Fig.4. And the detailed duration distribution of these videos is shown in Fig.5.
QA Pairs Annotation
When annotating the questions and answers for the collected videos, we devise two types of QA pair designs. Firstly, we formulate the QA pair as a binary classification task, where the question asks whether a certain event occurs in the given video, and the model is required to make a choice between “yes” and “no”. Besides, we also ask the model to freely describe the video contents in an open-ended manner. And then we match the detailed description with the ground-truth events.
Evaluation Metric Calculation
Binary Classification.
Since we have formulated the answer as either ”yes” or ”no”, we simply count the first word of the generated response and compare it with the ground-truth answer. The accuracy is provided for every category respectively.
Detailed Description Matching.
For each generated detailed description, we employ the powerful GPT-4o to examine whether hallucinations have occurred by comparing generated responses with ground-truth descriptions. More details about templates for prompting GPT-4o can be seen in the Appendix.
Figure 4: Video category and questions distribution. “Entire” implies that the entire video depicts the rare event. “Interleave” implies that common and rare events interleavedly occur in the video. “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
Figure 5: Video duration distribution
Temporal Contrastive Decoding
In this section, we propose Temporal Contrastive Decoding (TCD) to alleviate the event hallucination in VideoLLMs. We first explore the influence of sampled frame numbers on the accuracy and then introduce details about our proposed TCD method.
Impact of Sampled Frame Numbers
While current VideoLLMs exhibit a universal deficiency in dealing with event hallucination problems, their accuracy on our binary classification task can be influenced by the number of sampled frames. We believe that more sampled frames can enhance the temporal modeling of events, thereby improving the comprehension of VideoLLMs. To verify our hypothesis, we conduct the experiment to explore the overall accuracy of VideoLLMs on EventHallusion with different sampled frame numbers. As results have shown in Figure 6, the overall accuracy of LLaVA-Next (Li et al. 2024a) and Vila (Lin et al. 2024) increases when more sampled frames are provided within a certain range. On the contrary, fewer sampled frames can compromise temporal information and cause event hallucinations.
Figure 6: Overall accuracy on binary classification tasks with different sampled frame numbers.
Contrast with distorted temporality
The above analysis indicates that sampled frame numbers can influence the modeling of temporal information in VideoLLMs, which inspires us to adopt Contrastive Decoding (CD) (Li et al. 2022) to mitigate the event hallucination phenomenon. Given a textual input $x$ and the video $V$, the auto-regressive decoding process in VideoLLMs can be formulated as follows:
$p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)=\operatorname{softmax}\left(\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)\right)$ (1)
where $S(\cdot)$ represents sample operation on videos, $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$. $y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability in the softmax processed logit distribution.
To construct the temporal modeling distorted video for comparison, we use the sample operation to extract fewer frames, the original sampling operation can be formulated as follows:
$S(V,m)=\left\{I_{t_{1}},I_{t_{2}},\ldots,I_{t_{m}}\right\},$
$t_{i}=\left\lfloor\frac{(i-1)\cdot n}{m-1}\right\rfloor+1,\quad\text{for }i=1,2,\ldots,m$ (2)
where $n$ and $m$ represent the whole frames and sampled frames numbers, we perform the sample operation uniformly along the time. We sample fewer $m^{\prime}<m$ frames $S(V,m^{\prime})$ as the comparison for the contrastive decoding. The whole decoding process can be defined as:
$p_{tcd}$ $\left(y\mid S(V,m),R(S(V,m^{\prime})),x;\theta\right)=$
$\operatorname{softmax}\Big{[}(1+\alpha)\operatorname{logit}_{\theta}\left(y\mid S(V,m),x;\theta\right)$
$\quad-\alpha\operatorname{logit}_{\theta}\left(y\mid S(V,m^{\prime}),x;\theta\right)\Big{]}$ (3)
We adopt the hyperparameter $\alpha$ here to control the strength of contrastive decoding as utilized in (Leng et al. 2024).
As has been considered in (Li et al. 2022), the contrastive decoding process is performed for all tokens to be generated, which will impact the generation of non-visual-related content. Thus, we adopt the adaptive plausible constraint to control tokens for contrastive decoding:
$\begin{array}[]{l}\mathcal{V}_{\text{head }}\left(y_{<t}\right)=\left\{y_{t}\in\mathcal{V}:\right.\\ \left.p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)\geq\beta\max_{w}p\left(w\mid S(V),x,y_{<t};\theta\right)\right\}\end{array}$ (4)
Our TCD is only applied to tokens in the subset $\mathcal{V}_{\text{head }}$, where the hyperparameter $\beta$ in $[0,1]$ controls the truncation of logit distribution.
Table 1: Performance comparison of VideoLLMs on EnventHallusion. We exhibit the accuracy of both open-sourced and closed-sourced models on binary classification task and detailed description matching task on our three types of videos respectively. We use red to highlight the best results in close-source VideoLLMs and blue for open-source VideoLLMs. In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
Models Success Entire Interleave Misleading Overall
Binary Desc. Binary Desc. Binary Desc. Binary Desc.
Open-source Models
VideoChatGPT (Maaz et al. 2023) 100.00 14.91 5.50 79.80 3.63 21.57 34.74 47.19 11.59
VideoChat2 (Li et al. 2024b) 83.13 16.67 4.59 58.55 1.55 22.55 24.21 37.90 7.81
PLLaVA (Xu et al. 2024) 99.51 45.61 16.51 68.91 3.11 81.37 51.58 65.53 18.39
LLaMA-VID (Li, Wang, and Jia 2023) 100.00 30.70 16.51 98.96 7.77 43.14 43.16 66.01 18.64
LLaVA-NeXT-Video (Li et al. 2024a) 99.80 49.12 9.17 81.35 4.66 71.57 46.32 69.93 15.87
VILA (Lin et al. 2024) 100.00 53.51 20.18 62.18 17.62 83.33 80.00 65.04 33.25
ShareGPT4Video (Chen et al. 2024) 89.49 11.40 0.00 93.78 5.18 6.86 30.53 49.14 9.82
Video-LLaVA (Lin et al. 2023) 100.00 30.70 8.26 67.36 7.25 41.18 46.32 50.61 16.88
Close-source Models
GPT-4o (Achiam et al. 2023) 99.76 65.79 34.86 90.67 54.92 92.16 83.16 84.11 56.17
Gemini-1.5 (Team et al. 2023) 97.31 70.16 38.53 77.72 40.93 96.08 80.00 80.20 49.62
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We exhibit the accuracy of the binary classification task. The best results of each category are bolded.
Model Decoding Entire Interleave Misleading Overall
LLaVA-NeXT-Video (Li et al. 2024a) Default 49.12 81.35 71.57 69.93
TCD 50.88 84.46 74.51 72.62
VideoChat2 (Li et al. 2024b) Default 16.67 58.55 22.55 37.90
TCD 17.54 60.62 25.49 39.85
Vila (Lin et al. 2024) Default 53.51 62.18 83.33 65.04
TCD 50.88 70.98 84.31 68.70
Experiments
In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, incorporating both open-sourced and close-sourced MLLMs. Subsequently, we assess our TCD method on our EventHallusion with three VideoLLMs and demonstrate its effectiveness in improving the robustness of VideoLLMs against event hallucinations.
Evaluation on EventHallusion
Evaluated VideoLLMs.
We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT (Maaz et al. 2023), VideoChat2 (Li et al. 2024b), PLLaVA (Xu et al. 2024), LLaMA-VID (Li, Wang, and Jia 2023), VideoLaVIT (Jin et al. 2024), LLaVA-NeXT-Video (Li et al. 2024a), VILA (Lin et al. 2024), ShareGPT4Video (Chen et al. 2024), and Video-LLaVA (Lin et al. 2023). For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (Achiam et al. 2023), GPT-4o and Gemini-1.5 (Team et al. 2023), to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
For the inference setting of each VideoLLM, we follow their original setting including conversation mode and other hyperparameters. For the number of frames, we sample at least 12 frames for each VideoLLM to ensure the complete preservation of event information. In addition, to ensure the stability and reproduction of our results, we adopt the greedy search as the default decoding strategy for all open-sourced models. For the close-sourced models, we sample 8 frames.
Main Results
We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As results have shown, open-sourced VideoLLMs manifest a universal weakness in event hallucinations. In contrast, close-sourced VideoLLMs are more robust against event hallucinations across all categories. We delve into a deeper analysis as follows.
For the evaluation of the sensitivity of VideoLLMs towards video content, the performance on the detailed description matching of Interleave is much lower than the result on the binary classification task, which suggests that VideoLLMs may follow the question as the cue to search rare events. However, they are prone to overlook rare events in their detailed descriptions when no cues are provided. Meanwhile, VideoLLMs have difficulty in processing both binary classification and detailed description matching on Interleave. This indicates the over-reliance of VideoLLMs on language priors. Notably, VideoLLMs show worse capacity on the description matching of Interleave than Entire, which proves that common events can strongly induce VideoLLMs to generate hallucinations than with only rare events.In Misleading, despite their high accuracy on description matching, results on binary classification indicate that VideoLLMs can be affected by questions and generate event hallucinations.
Figure 7: Accuracy of TCD on binary classification of EventHallusion with different numbers of sampled frames in the distorted video. Base indicates the results of standard decoding with 32 frames sampled.
Evaluating TCD on EventHallusion
VideoLLM Baselines.
To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (Li et al. 2024a), VideoChat2 (Li et al. 2024b), and Vila (Lin et al. 2024) as the foundation backbones. LLaVA-NeXT-Video extends the ability of LLaVA (Liu et al. 2024a) based on the projection layer design, while VideoChat2 adopts the Q-Former (Li et al. 2023a) to bridge the gap between videos and LLM. Vila, on the other hand, employs enhanced visual-language pre-training for better video comprehension. Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
Implementation Details.
We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for Vila, and 16 and 4 frames for VideoChat2. We follow (Leng et al. 2024) and set $\alpha=1$, $\beta=0.1$. We use greedy search as the basic decoding paradigm.
Main Results
We exhibit the results of the binary classification task on EventHallusion in Tab. 2. As shown in Tab. 2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms. However, while TCD shows consistent and notable improvements on Interleave and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed. This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
Effect of Sampled Frames in Distorted Video.
We further explore the effect of sampled frames in distorted video on our TCD. Specifically, we adopt LLaVA-NeXT-Video as the backbone model with 32 frames sampled for the original video input and adjust the numbers of sampled frames in the distorted video. We conduct this experiment on the binary classification task of EventHallusion. As results have shown in Figure 7, TCD maintains a steady improvement in the range of 4 to 16 sampled frames. However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates. We attribute this phenomenon to the excessive destruction of the temporal modeling in the VideoLLM with too few frames incorporated. This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
Conclusion
In this paper, we systematically investigated the event-related hallucination problem of existing VideoLLMs. To provide a reliable evaluation benchmark to inspect such hallucination degree of VideoLLMs, we curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves. In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems of existing VideoLLMs in a lightweight plug-and-play manner. We hope our research can benefit further research in the field of reducing the hallucinations of VideoLLMs.
Appendix A A. Automatic Evaluation For Detailed Description Matching
We apply GPT-4o to evaluate the detailed description matching task of our Entire, Interleave, and Misleading categories respectively. For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event. And for Interleave, we prompt GPT4o to focus on the occurrence of the interleaved rare event. Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events. In addition, we set up the default system prompt to better guide GPT-4o’s evaluation, which is shown in Figure 8. The detailed prompts are shown in Figure  9, 10, 11, respectively.
Figure 8: System prompt used in all automatic evaluation.
Appendix B B. Qualitative Examples
Data examples of EventHallusion.
We present examples of Entire, Interleave, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs. As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
Examples of GPT-4o’s Evaluation.
We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Interleave, and Misleading respectively in Figure 15, 16, 17. As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
Examples of TCD.
We present examples of our TCD in mitigating video event hallucinations. Specifically, we exhibit the generated response with and without TCD for LLaVA-NeXT-Video and VideoChat2 on our EventHallusion in Figure 18 and 19.
Figure 9: Prompt used for detailed description matching in Entire.
Figure 10: Prompt used for detailed description matching in Interleave.
Figure 11: Prompt used for detailed description matching in Misleading.
Figure 12: Examples of Entire with responses of VideoLLMs.
Figure 13: Examples of Interleave with responses of VideoLLMs.
Figure 14: Examples of Misleading with responses of VideoLLMs.
Figure 15: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Entire.
Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Interleave.
Figure 17: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Misleading.
Figure 18: Examples of TCD with LLaVA-NeXT-Video on our EventHallusion.
Figure 19: Examples of TCD with VideoChat2 on our EventHallusion.

### v2

@@
 Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain.
-To mitigate this gap, we first propose EventHallusion, a novel benchmark that focuses on assessing the VideoLMMs’ hallucination phenomenon on video event comprehension.
-Based on the observation that existing VideoLLMs are entangled with the priors stemming from their foundation models, our EventHallusion is curated by meticulously collecting videos and annotating questions to intentionally mislead the VideoLLMs into interpreting events based on these priors rather than accurately understanding the video content.
+To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis.
+From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases.
 On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs.
-The proposed TCD suppresses the model’s preference toward their priors by comparing the original video with a constructed counterpart, whose temporal cues are disrupted, during the autoregressive decoding stage.
-Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we find that the open-source models suffer significantly from hallucination problems, whereas the closed-source models perform markedly better.
-By further equipping open-sourced VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark.
+The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are disrupted.
+Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better.
+By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark.
 Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
-Introduction
-The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) (Brown et al.
-2020; Touvron et al.
-2023; Dettmers et al.
-2024) and Multimodal Large Language Models (MLLMs) (Alayrac et al.
-2022; Liu et al.
-2024a; Bai et al.
-2023; Jiao et al.
-2024b, a).
-Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  (Jiao et al.
-2024a; Yin et al.
-2023b).
-Despite their great success, they still face hallucination problems, where models generate plausible yet factually incorrect responses (Leng et al.
-2024), potentially threatening reliability and security in real-world applications.
+Figure 1: Illustration of two types of underlying reasons that may lead to existing VideoLLMs’ hallucinations.
+We use red to indicate the hallucinations in answers generated by existing state-of-the-art VideoLLMs, namely Video-LLaVA and ShareGPT4Video.
+1 Introduction
+The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 36, 10] and Multimodal Large Language Models (MLLMs) [4, 28, 5, 15, 13].
+Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [13, 41].
+Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [18], potentially threatening reliability and security in applying them in real-world scenarios [33, 14].
 To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.
-(Li et al.
-2023b; Fu et al.
-2023; Liu et al.
-2023b).
-However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  (Ma et al.
-2023; Li, Wang, and Jia 2023).
-To mitigate this gap, two recent novel benchmarks (Wang et al.
-2024; Guan et al.
-2024) have been proposed.
-Specifically, as illustrated in Fig.1, HallusionBench (Guan et al.
-2024) reverses all frames within a video and then queries the model to determine whether the event described in the original captions occurs in the altered video.
-VideoHallucer (Wang et al.
-2024) employs long videos containing multiple events and then queries the model to assess the chronological order of these events.
-Generally, compared to those focused on hallucinations in static images, benchmarks that assess event hallucinations in dynamic videos lag behind.
-When delving into VideoLLMs’ hallucinations on events, we found that the major causes can be attributed to language priors of LLMs (Agrawal, Batra, and Parikh 2016; Agarwal, Shetty, and Fritz 2020) and statistical bias of the inherited ImageLLM’s training corpus (Leng et al.
-2024).
-As shown in Fig.1, when presented with a woman and a bike, Video-LLaVA (Lin et al.
-2023), LLaVA-NeXT-Video (Li et al.
-2024a), and ShareGPT4Video (Chen et al.
-2024) associate them with the intuitive assumption of ”riding the bicycle”, rather than recognizing the clear fact of ”carrying the bicycle”.
-However, this essential event-related hallucination cannot be effectively assessed in both VideoHallucer (Wang et al.
-2024) and HallusionBench (Guan et al.
-2024).
-As shown in Fig.1, VideoHallucer has already provided the correct events in the question, merely requiring the model to determine their chronological orders without analyzing the events themselves.
-Although HallusionBench addresses event-oriented hallucinations, the frame reversal operations employed to create test samples may deviate from the natural distribution of videos.
-Figure 1: Comparison of our proposed EventHallusion with previous benchmarks.
-Different from requiring models to locate events or comprehend reversed videos, our EventHallusion focuses on evaluating inherent priors-oriented hallucinations concerning events in VideoLLMs.
-Green and red represent the original and hallucination settings for evaluation in HallusionBench and VideoHallucer.
-At the bottom, we exhibit an example in our EventHallusion where Video-LLaVA, LLaVA-NeXT-Video, and ShareGPT4Video generate hallucinations towards the video event with hallucinated content marked by red.
-Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates state-of-the-art VideoLLMs for hallucinations concerning events that are susceptible to the model’s inherent priors.
-The primary challenge in developing the EventHallusion is sourcing appropriate video data.
-Since existing publicly available video datasets, such as AcitivityNet (Caba Heilbron et al.
-2015), randomly sample video fragments from daily life.
-As a result, their data distribution naturally reflects the priors resulting from the high-frequency co-occurrence of objects and event bindings (e.g., a person riding a bike in most cases).
-Therefore, instead of refactoring samples from existing video datasets, we manually collect videos and annotate question-answer pairs which may lead VideoLLMs to interpret events based on their pre-existing priors from prevalent internet video platforms.
-Specifically, to comprehensively probe existing VideoLLMMs’ sensitivity to their learned priors, we categorize our benchmark into two basic categories as shown in Fig.2: (1) In the first category, all videos are meticulously selected to ensure that the included events occur infrequently.
-Within this category, we further split videos into two groups according to their event composition.
-Videos in the first group (i.e., “entire rare events” in Fig.2) depict a rare event throughout the entire duration.
-And videos in the second group (i.e., “interleaved common-rare events” in Fig.2) include a period depicting a common event, followed by a rare event unexpectedly happens.
-Compared with those in the first group, videos in the second group are more challenging because the initially presented common event acts as a demonstration to consolidate the model’s tendency to their priors, which is also proved in Tab.
-1.
-(2) In the second category, all questions are customized to include a plausible but incorrect common event as a distraction, and videos depict events that occur frequently.
-Instead of evaluating VideoLLM’s sensitivity to video content as in the first category, the testing samples in the second category are designed to assess the model’s sensitivity to question content, where the included misleading event descriptions can trigger the model to rely on its pre-existing priors.
-Through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination problems on our benchmark, which can provide valuable insights for further research.
-Moreover, inspired by a hallucination elimination approach for image LLMs (i.e., VCD (Leng et al.
-2024)), we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs.
-As previously analyzed, VideoLLMs can be easily influenced by their prior knowledge.
-Therefore, our TCD is designed to mitigate such biases and enhance the model’s objective judgments in event comprehension.
-Technically, for each given video, we first construct its counterpart with temporal downsampling and shuffling data augmentation operations.
-Both the original video and its counterpart are then fed into the VideoLLMs simultaneously to generate a pair of predicted logits, which are used for contrastive decoding similar to the approach in VCD (Leng et al.
-2024).
-In the proposed TCD, the created counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s prior knowledge.
-Subtracting these amplified priors in the logits space during the contrastive decoding process helps guide the model toward more reliable judgments.
+[22, 11, 29].
+Specifically, POPE examines dataset distributions to investigate the hallucination phenomenon in MLLMs, focusing on frequently occurring objects and common object combinations, while MME and MMBench assess hallucinations through systematic evaluations of object existence, attributes, and relations.
+However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  [31, 23].
+Two recent benchmarks [37, 12] further extend their evaluation to dynamic scenes with event-related queries.
+In pursuit of the rigorous evaluation, both of them curate more challenging event comprehension tasks.
+Specifically, HallusionBench [12] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video.
+VideoHallucer [37] selects long videos containing multiple events and then queries the model to evaluate the chronological order of these events.
+Although advancing the evaluation of event hallucination, they still suffer from the following limitations.
+On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse.
+On the other hand, merely judging the chronological order of events given in the questions fails to capture the model’s authentic ability to comprehend the detailed content of the events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [12] and VideoHallucer [37] is provided in the supplementary materials due to page limitations..
+When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs.
+Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge.
+As illustrated in the left column of Fig.1, both Video-LLaVA and ShareGPT4Video are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question.
+Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets.
+However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations.
+As shown in the right column of Fig.1, both Video-LLaVA and ShareGPT4Video intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
+Toward this end, in this paper, we propose EventHallusion, a novel benchmark that systematically evaluates event-related hallucinations of state-of-the-art VideoLLMs.
+From a hallucination attribution perspective, our EventHallusion is curated to assess the susceptibility of a VideoLLM to language priors and vision-language correlation biases, respectively.
+Specifically, our benchmark consists of two basic categories as shown in Fig.2: (1) Susceptibility to language priors: In this category, all questions are annotated with a misleading event designed to prompt the model to make decisions based on its language priors.
+Within such events, while objects align with those depicted in the video, their actions are deliberately misrepresented.
+(2) Susceptibility to vision-language correlation biases: In this category, all videos are meticulously selected to ensure that the included events are infrequent yet plausible.
+When presented with these infrequent events, the model can easily be misled to interpret the event as one that frequently co-occurs with the objects or scenes depicted in the video, based on its memory.
+It is worth mentioning that through comprehensive evaluation, we find that existing open-source VideoLLMs suffer from severe event hallucination issues on our benchmark, which could offer promising avenues for enhancing the VideoLLM’s capabilities in the future research.
+Figure 2: Illustration of examples in our EventHallusion benchmark.
+Questions are basically divided into two categories: Susceptibility to Language Prior and Susceptibility to Vision-Language Correlation Bias.
+The contents that induces VideoLLMs to generate event-related hallucinations are marked by red.
+Moreover, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs.
+Without any additional training budgets, our TCD rectifies the model’s bias to its prior during the decoding stage.
+Technically, for each given video, we first apply the temporal downsampling operation to create a modified counterpart.
+Both the original video and its modified counterpart are then input into the VideoLLMs.
+During the decoding stage, the logits prediction of the original video is adjusted by the logits prediction of the counterpart.
+The rectified logits are then used to sample the next token and continue the decoding process.
+Within the proposed TCD, the modified counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s event-related prior knowledge.
+Subtracting these amplified priors in the logits space during the decoding process helps guide the model to make more reliable judgments.
 As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
-In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We conduct comprehensive ablation studies to provide more in-depth analysis.
-Figure 2: Illustration of evaluation categories.
-Our EventHallusion is basically divided into two categories: Sensitivity to Video Content and Sensitivity to Question Content.
-The content that induces VideoLLMs to generate event-related hallucinations is marked by red.
-Figure 3: Framework of Temporal Contrastive Decoding.
-We construct the temporal modeling distorted video as the comparison group with fewer frames extracted and then contrast their logits to mitigate the event hallucination.
-Related Works
+In summary, our contributions are threefold: (1) We propose EventHallusion, a novel video QA benchmark focused on assessing event-related hallucinations in VideoLLMs; (2) We introduce Temporal Contrastive Decoding (TCD), a simple yet effective approach that can reduce hallucinations across a wide range of VideoLLMs in a plug-and-play manner; (3) We also conduct comprehensive ablation studies to provide more in-depth analysis.
+2 Related Works
 Video Large Language Models.
@@
 To enable MLLMs to comprehend video content and the concept of temporality, numerous methods have been put forward.
-Video-ChatGPT (Maaz et al.
-2023) utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA (Zhang, Li, and Bing 2023) introduces a Video Q-Former to aggregate frame-level representations.
-Vista-LLaMA (Ma et al.
-2023) preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality.
-LLaMA-VID (Li, Wang, and Jia 2023) proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information.
+Video-ChatGPT [32] utilizes spatial and temporal pooling operations to conduct video modeling, while Video-LLaMA [43] introduces a Video Q-Former to aggregate frame-level representations.
+Vista-LLaMA [31] preserves the equal distance between all visual tokens and any language tokens and proposes the temporal Q-Former projector to further enhance the representations of the video and the modeling of temporality.
+LLaMA-VID [23] proposes a dual-token paradigm, using a context token and a content token to represent each frame, which effectively supports the comprehension of long videos and maintains temporal information.
 However, despite their excellent performance on existing benchmarks, these methods still suffer the hallucination problem in temporality.
@@
 Despite the extraordinary advancement of MLLMs, it is of great significance to conduct quantitative evaluations of existing models.
-To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made (Fu et al.
-2023; Liu et al.
-2023b; Li et al.
-2024c).
+To thoroughly evaluate the performance of image MLLMs, a wide scope of efforts has been made [11, 29, 24].
 Their evaluation comprises various aspects, including perception, reasoning, and cognition.
 With the development of VideoLLMs, several benchmarks have also been proposed to assess video comprehension capacities.
-Tempcompass (Liu et al.
-2024b) focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order.
-FunQA (Xie et al.
-2023) proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
+Tempcompass [30] focuses on the thorough evaluation of the temporal perception ability of VideoLLMs, including action, speed, direction, attribute change, and event order.
+FunQA [38] proposes utilizing humorous, creative, and magic videos to assess the in-depth comprehension ability of existing VideoLLMs.
 Despite previous works sharing similarities with ours, there exist several critical differences.
@@
 To alleviate the impact of hallucinations in image MLLMs, several methods have been proposed.
-Woodpecker (Yin et al.
-2023a) adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano (Lee et al.
-2023) proposes a critique-revise-decide process to self-revise their initial generated answers.
-Besides, LRV-Instruction (Liu et al.
-2023a) and HalluciDoctor (Yu et al.
-2024) propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
+Woodpecker [40] adopts a post-hoc correction mechanism that utilizes additional extracted visual cues to rectify previously generated responses, while Volcano [17] proposes a critique-revise-decide process to self-revise their initial generated answers.
+Besides, LRV-Instruction [27] and HalluciDoctor [42] propose to mitigate this problem from the data perspective to provide more balanced, high-quality instruction tuning datasets.
 Some other works propose to alleviate hallucination during the inference phase, in which the contrastive decoding (CD) method has been widely adopted due to its excellent performance and the absence of additional training.
-VCD (Leng et al.
-2024) employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD (Zhu et al.
-2024) constructs the comparison pair by introducing the image-biased and original MLLMs.
+VCD [18] employs the original and distorted images for comparison in contrastive decoding, thereby reducing the impact of statistical bias and language priors, while IBD [44] constructs the comparison pair by introducing the image-biased and original MLLMs.
 Despite the extensive exploration of adopting CD in ImageLLMs, there is still a lack of effort in employing CD to handle hallucinations in VideoLLMs.
 Inspired by this, we propose Temporal Contrastive Decoding (TCD) to mitigate event hallucinations in VideoLLMs.
-EventHallusion Benchmark
-In this section, we first analyze the vulnerabilities of VideoLLMs emphasized by each category in our EventHallusion benchmark.
-Then, we elaborate on the detailed curation process of video data and question-answer pair annotation.
-Finally, we introduce the metrics calculation approach.
-Vulnerability Dimension Analysis
-To comprehensively assess a VideoLLM’s sensitivity to its priors across diverse aspects, we employ two categories to examine the influence of video and question content, respectively.
-Sensitivity to Video Content
-In this category, we intentionally collect videos depicting events that infrequently occur in daily life (e.g., a person carrying a bike, as shown in Fig.1).
-Such videos can easily trigger the prior knowledge learned from foundational ImageLLMs (noting that existing open-source VideoLLMs are all derived from certain ImageLLMs).
-As a result, the VideoLLM may rely on these strong visual priors rather than accurately interpreting the temporal cues of the video.
-Sensitivity to Question Content
-In this category, we select videos depicting events that commonly occur in daily life to exclude the influence of the visual priors mentioned previously.
-Additionally, we introduce a frequently occurring event that is inconsistent with the ground truth depicted in the video (e.g., “eating” in the Fig.2).
-Such customized questions can easily trigger priors of the LLM which obtains knowledge from enormous textual corpus.
-As a result, the VideoLLM may rely on its language priors to find a shortcut in answering the question.
-Video Data Collection
-When collecting video sources for our EventHallusion, we adhere to the following two rules to ensure high quality.
-Firstly, the objects and inter-object interaction details should be presented clearly to avoid visual detail missing.
-Secondly, the scene presented in the video is kept consistent without any shot cuts to avoid the adverse effects of environmental variation.
-In total, we collect 400 eligible videos from multiple open video platforms on the Internet.
-To avoid data leakage, we have blurred the captions in the video, thereby encouraging models to analyze video events solely through visual content.
-These 400 videos cover 7 domains including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, and Nature as shown in Fig.4.
-And the detailed duration distribution of these videos is shown in Fig.5.
-QA Pairs Annotation
-When annotating the questions and answers for the collected videos, we devise two types of QA pair designs.
-Firstly, we formulate the QA pair as a binary classification task, where the question asks whether a certain event occurs in the given video, and the model is required to make a choice between “yes” and “no”.
-Besides, we also ask the model to freely describe the video contents in an open-ended manner.
-And then we match the detailed description with the ground-truth events.
-Evaluation Metric Calculation
-Binary Classification.
-Since we have formulated the answer as either ”yes” or ”no”, we simply count the first word of the generated response and compare it with the ground-truth answer.
-The accuracy is provided for every category respectively.
-Detailed Description Matching.
-For each generated detailed description, we employ the powerful GPT-4o to examine whether hallucinations have occurred by comparing generated responses with ground-truth descriptions.
-More details about templates for prompting GPT-4o can be seen in the Appendix.
+3 EventHallusion Benchmark
+In this section, we first elaborate on the data curation process of two basic categories of our EventHallusion benchmark.
+Afterward, we demonstrate the distribution statistics of the whole EventHallusion benchmark.
+Finally, we introduce the specific evaluation protocol of different question types.
+3.1 Data Curation
+Susceptibility to language priors.
+In this category, we first collect videos describing the events that occur in daily life from the internet.
+To avoid information leakage, we blur the text captions printed on videos, thereby encouraging models to analyze events solely based on visual content.
+However, when annotating the questions, we explicitly include another frequently occurring event that is inconsistent with the video content to trigger the language priors within the LLM.
+With these “misleading triggers”, we formulate questions as a binary choice problem as shown below:
+“Did the [subject] [event] happen in the video?”
+Here, the combination of [subject] and [event] correspond to the “misleading triggers” aforementioned. In Fig.2, we refer to this type of sample as “Common Events with Misleading Questions”.
+Susceptibility to vision-language correlation biases.
+In this category, the primary challenge lies in sourcing the appropriate video data.
+Since existing publicly available video datasets (e.g, ActivityNet [7], [9], [34]) collect videos from daily life, the events they include cater to the vision-language correlation biases stored in VideoLLMs, resulting in high recognition precision.
+Therefore, instead of refactoring existing video datasets, we meticulously collect videos from prevalent Internet video platforms and filter them with rigorous cross-validation.
+Specifically, we begin by manually collecting videos that depict infrequent events based on human consensus.
+To minimize personal bias, we engage three annotators to assess whether the events in these selected videos are indeed uncommon.
+Only videos that receive unanimous positive judgments from all annotators are retained as testing samples.
+Similar to the first category, we also employ the blurring operation on videos to prevent information leakage.
+We further divide these videos into two groups: (1) Videos where infrequent events occur throughout the entire duration, referred to as “Entire Rare Events” in Fig.2.
+(2) Videos containing a mix of both frequent and infrequent events, referred to as “Mix Common-Rare Events” in Fig.2.
+For constructing questions, we design both binary and open-ended types for comprehensive evaluation.
+Specifically, the binary questions ask the model whether a plausible but incorrect common event occurs in the given video.
+The open-ended questions, on the other hand, simply prompt the model to freely describe the video content.
+Formally, binary and open-ended questions for the “Entire” category can be formulated as presented below:
+Binary: “Did the [subject] [event] in the video?”,
+Open-Ended: “Please describe this video in detail.”,
+For the “Mix” category, since both common and rare events can happen in the video, we reformulate the binary questions as below:
+“Did the [subject] [event] throughout the entire video?”,
+“Did any accident or anything unexpected happen in the video?”,
+Dataset distribution statistics.
+In total, our EventHallusion benchmark comprises 400 videos and 711 questions.
+To ensure comprehensive coverage across diverse domains, these videos span seven scenarios, including Pet&Animal, Sports Competition, Food&Drink, Gym Exercises, Vehicle, Life Record, Nature, as shown in Fig.3.
+The distribution of video lengths in our benchmark is presented in Fig.4, showing that most videos are relatively short.
+This is because longer videos often involve significant scene fluctuations, which introduce additional challenges for VideoLLMs beyond the uni-modal and multi-modal biases we aim to examine in this paper.
 Figure 3: Video category and questions distribution.
 “Entire” implies that the entire video depicts the rare event.
-“Interleave” implies that common and rare events interleavedly occur in the video.
+“Mix” implies that common and rare events interleavedly occur in the video.
 “Misleading” implies the video with the common event, but with its question customized to include misleading event demonstration.
-Figure 5: Video duration distribution
-Temporal Contrastive Decoding
-In this section, we propose Temporal Contrastive Decoding (TCD) to alleviate the event hallucination in VideoLLMs.
-We first explore the influence of sampled frame numbers on the accuracy and then introduce details about our proposed TCD method.
-Impact of Sampled Frame Numbers
-While current VideoLLMs exhibit a universal deficiency in dealing with event hallucination problems, their accuracy on our binary classification task can be influenced by the number of sampled frames.
-We believe that more sampled frames can enhance the temporal modeling of events, thereby improving the comprehension of VideoLLMs.
-To verify our hypothesis, we conduct the experiment to explore the overall accuracy of VideoLLMs on EventHallusion with different sampled frame numbers.
-As results have shown in Figure 6, the overall accuracy of LLaVA-Next (Li et al.
-2024a) and Vila (Lin et al.
-2024) increases when more sampled frames are provided within a certain range.
-On the contrary, fewer sampled frames can compromise temporal information and cause event hallucinations.
-Figure 6: Overall accuracy on binary classification tasks with different sampled frame numbers.
-Contrast with distorted temporality
-The above analysis indicates that sampled frame numbers can influence the modeling of temporal information in VideoLLMs, which inspires us to adopt Contrastive Decoding (CD) (Li et al.
-2022) to mitigate the event hallucination phenomenon.
-Given a textual input $x$ and the video $V$, the auto-regressive decoding process in VideoLLMs can be formulated as follows:
-$p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)=\operatorname{softmax}\left(\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)\right)$ (1)
-where $S(\cdot)$ represents sample operation on videos, $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$.
-$y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability in the softmax processed logit distribution.
-To construct the temporal modeling distorted video for comparison, we use the sample operation to extract fewer frames, the original sampling operation can be formulated as follows:
-$S(V,m)=\left\{I_{t_{1}},I_{t_{2}},\ldots,I_{t_{m}}\right\},$
-$t_{i}=\left\lfloor\frac{(i-1)\cdot n}{m-1}\right\rfloor+1,\quad\text{for }i=1,2,\ldots,m$ (2)
-where $n$ and $m$ represent the whole frames and sampled frames numbers, we perform the sample operation uniformly along the time.
-We sample fewer $m^{\prime}<m$ frames $S(V,m^{\prime})$ as the comparison for the contrastive decoding.
-The whole decoding process can be defined as:
-$p_{tcd}$ $\left(y\mid S(V,m),R(S(V,m^{\prime})),x;\theta\right)=$
-$\operatorname{softmax}\Big{[}(1+\alpha)\operatorname{logit}_{\theta}\left(y\mid S(V,m),x;\theta\right)$
-$\quad-\alpha\operatorname{logit}_{\theta}\left(y\mid S(V,m^{\prime}),x;\theta\right)\Big{]}$ (3)
-We adopt the hyperparameter $\alpha$ here to control the strength of contrastive decoding as utilized in (Leng et al. 2024).
-As has been considered in (Li et al.
-2022), the contrastive decoding process is performed for all tokens to be generated, which will impact the generation of non-visual-related content.
-Thus, we adopt the adaptive plausible constraint to control tokens for contrastive decoding:
-$\begin{array}[]{l}\mathcal{V}_{\text{head }}\left(y_{<t}\right)=\left\{y_{t}\in\mathcal{V}:\right.\\ \left.p\left(y_{t}\mid S(V),x,y_{<t};\theta\right)\geq\beta\max_{w}p\left(w\mid S(V),x,y_{<t};\theta\right)\right\}\end{array}$ (4)
-Our TCD is only applied to tokens in the subset $\mathcal{V}_{\text{head }}$, where the hyperparameter $\beta$ in $[0,1]$ controls the truncation of logit distribution.
+Figure 4: The distribution of video length within our EventHallusion benchmark.
+3.2 Evaluation Protocol
+Binary question answering.
+Following previous benchmarks [37, 22], we compare the first word of the model’s answer with the ground-truth answer to determine whether the model answered the question correctly.
+This approach is justified by the VideoLLM’s exceptional instruction-following capability, which typically enables it to answer binary choice questions with “yes” or “no” at the beginning.
+We calculate the accuracy as the evaluation metric for binary questions.
+Open-ended question answering.
+For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o, using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [30].
+We also use accuracy as the evaluation metric.
+Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
+Figure 5: Framework of Temporal Contrastive Decoding.
+We construct the temporal modeling distorted video as the comparison group with fewer frames extracted and then contrast their logits to mitigate the event hallucination.
+4 Temporal Contrastive Decoding
+To reduce the aforementioned hallucinations in VideoLLMs, the most straightforward approach is to further enrich debiased data for finetuning.
+However, this approach is particularly resource-intensive for collecting such videos and annotating QA pairs.
+To address this, we present our Temporal Contrastive Decoding (TCD) method, which effectively alleviates hallucinations in videos without requiring additional training efforts.
+Motivated by the previous work [18], our TCD method aims to reduce event hallucinations by adjusting the distribution of predicted logits at each decoding step, using a contrastive counterpart that is more aligned with the model’s inherent biases.
+For clarity, we first review the standard auto-regressive decoding process of a VideoLLM.
+Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
+$z_{t}^{ori}=\operatorname{logit}_{\theta}\left(y_{t}\mid V,x,y_{<t}\right)$ (1)
+where $y_{t}$ is the token to be generated at time step $t$, and $y_{<t}$ denotes the generated token sequence before time step $t$.
+$y_{t}$ is based on $x$, $V$, and all previously generated tokens $y_{<t}$ by selecting the token with the highest probability.
+When constructing a contrastive counterpart, we chronologically downsample the original video as shown in the Fig.5, aiming to weaken the temporal cues while emphasizing object cues.
+Therefore, the contrastive counterpart’s probability prediction $z_{t}^{con}$ at timestep $t$ can be formulated as:
+$z_{t}^{con}=\operatorname{logit}_{\theta}\left(y_{t}\mid S(V),x,y_{<t}\right)$ (2)
+where $S(\cdot)$ represents the temporal downsampling operation.
+With the generated $z_{t}^{ori}$ and $z_{t}^{con}$, the modulated probability distribution $p_{t}$ at the current timestep can be calculated as:
+$p_{t}=\operatorname{softmax}\Big{[}(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}\Big{]}$ (3)
+where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding.
+However, applying such a contrastive process throughout the entire decoding stage would lead to significant model degradation, as it is not always consistent for $z_{t}^{ori}$ and $z_{t}^{con}$ to have mismatched distribution.
+When $z_{t}^{ori}$ and $z_{t}^{con}$ are consistent and correct, employing Eq.3 to modulate the probability impairs the model’s judgment.
+Therefore, we introduce another hyperparameter $\beta$ to constrain the token set used for TCD.
+The motivation behind this design is twofold: when the model exhibits high confidence in a specific token, we conduct TCD on a smaller subset of tokens with similarly high confidence.
+Conversely, when the model’s confidence is distributed more evenly across tokens, we apply a lower threshold, enabling TCD on a broader range of tokens.
 Table 1: Performance comparison of VideoLLMs on EnventHallusion.
@@
 In addition, we show the overall success rate of each model on our binary classification task that answers ”yes” or ”no” in the first word.
-Models Success Entire Interleave Misleading Overall
-Binary Desc. Binary Desc. Binary Desc. Binary Desc.
+Models Entire Mix Misleading Overall
+Binary Desc. Binary Desc. Binary Binary Desc.
 Open-source Models
-VideoChatGPT (Maaz et al. 2023) 100.00 14.91 5.50 79.80 3.63 21.57 34.74 47.19 11.59
-VideoChat2 (Li et al. 2024b) 83.13 16.67 4.59 58.55 1.55 22.55 24.21 37.90 7.81
-PLLaVA (Xu et al. 2024) 99.51 45.61 16.51 68.91 3.11 81.37 51.58 65.53 18.39
-LLaMA-VID (Li, Wang, and Jia 2023) 100.00 30.70 16.51 98.96 7.77 43.14 43.16 66.01 18.64
-LLaVA-NeXT-Video (Li et al. 2024a) 99.80 49.12 9.17 81.35 4.66 71.57 46.32 69.93 15.87
-VILA (Lin et al. 2024) 100.00 53.51 20.18 62.18 17.62 83.33 80.00 65.04 33.25
-ShareGPT4Video (Chen et al. 2024) 89.49 11.40 0.00 93.78 5.18 6.86 30.53 49.14 9.82
-Video-LLaVA (Lin et al. 2023) 100.00 30.70 8.26 67.36 7.25 41.18 46.32 50.61 16.88
+VideoChatGPT [32] 14.91 5.50 56.99 3.63 21.57 36.43 4.30
+VideoChat2 [21] 16.67 4.59 12.44 1.55 22.55 16.14 2.64
+PLLaVA [39] 45.61 16.51 58.55 3.11 81.37 60.64 6.05
+LLaMA-VID [23] 30.70 16.51 73.58 7.77 43.14 54.03 10.90
+LLaVA-NeXT-Video [19] 49.12 9.17 70.47 4.66 71.57 64.79 6.29
+VILA [26] 53.51 20.18 58.55 17.62 83.33 63.33 18.54
+ShareGPT4Video [8] 11.40 0.00 67.88 5.18 6.86 49.14 9.82
+Video-LLaVA [25] 30.70 8.26 57.51 7.25 41.18 45.97 7.62
 Close-source Models
-GPT-4o (Achiam et al. 2023) 99.76 65.79 34.86 90.67 54.92 92.16 83.16 84.11 56.17
-Gemini-1.5 (Team et al. 2023) 97.31 70.16 38.53 77.72 40.93 96.08 80.00 80.20 49.62
+GPT-4o [1] 65.79 34.86 83.67 54.92 92.16 81.11 47.69
+Gemini-1.5 [35] 70.16 38.53 74.72 40.93 96.08 78.20 40.07
 Table 2: Results of TCD on EventHallusion.
@@
 The best results of each category are bolded.
-Model Decoding Entire Interleave Misleading Overall
-LLaVA-NeXT-Video (Li et al. 2024a) Default 49.12 81.35 71.57 69.93
-TCD 50.88 84.46 74.51 72.62
-VideoChat2 (Li et al. 2024b) Default 16.67 58.55 22.55 37.90
-TCD 17.54 60.62 25.49 39.85
-Vila (Lin et al. 2024) Default 53.51 62.18 83.33 65.04
-TCD 50.88 70.98 84.31 68.70
-Experiments
-In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, incorporating both open-sourced and close-sourced MLLMs.
-Subsequently, we assess our TCD method on our EventHallusion with three VideoLLMs and demonstrate its effectiveness in improving the robustness of VideoLLMs against event hallucinations.
-Evaluation on EventHallusion
-Evaluated VideoLLMs.
+Model Decoding Entire Mix Misleading Overall
+LLaVA-NeXT-Video [19] Default 49.12 70.47 71.57 64.79
+TCD 50.88 73.58 74.51 67.48
+VideoChat2 [21] Default 16.67 47.67 22.55 32.76
+TCD 17.54 49.74 25.49 34.72
+VILA [26] Default 53.51 58.55 83.33 63.33
+TCD 50.88 66.32 84.31 66.50
+5 Experiments
+In this section, we first evaluate multiple prevalent VideoLLMs on our EventHallusion, including both open-sourced and private MLLMs.
+Subsequently, we examine the effectiveness of our proposed TCD, and also discuss its more specific design choices.
+5.1 Evaluation on EventHallusion
+5.1.1 Selected VideoLLMs for Evaluation
 We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion.
-Open-sourced VideoLLMs comprise VideoChatGPT (Maaz et al.
-2023), VideoChat2 (Li et al.
-2024b), PLLaVA (Xu et al.
-2024), LLaMA-VID (Li, Wang, and Jia 2023), VideoLaVIT (Jin et al.
-2024), LLaVA-NeXT-Video (Li et al.
-2024a), VILA (Lin et al.
-2024), ShareGPT4Video (Chen et al.
-2024), and Video-LLaVA (Lin et al.
-2023).
-For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V (Achiam et al.
-2023), GPT-4o and Gemini-1.5 (Team et al.
-2023), to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
+Open-sourced VideoLLMs comprise VideoChatGPT [32], VideoChat2 [21], PLLaVA [39], LLaMA-VID [23], VideoLaVIT [16], LLaVA-NeXT-Video [19], VILA [26], ShareGPT4Video [8], and Video-LLaVA [25].
+For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o and Gemini-1.5 [35], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
 For the inference setting of each VideoLLM, we follow their original setting including conversation mode and other hyperparameters.
@@
 For the close-sourced models, we sample 8 frames.
-Main Results
+5.1.2 Performance Analysis
 We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task.
-As results have shown, open-sourced VideoLLMs manifest a universal weakness in event hallucinations.
-In contrast, close-sourced VideoLLMs are more robust against event hallucinations across all categories.
-We delve into a deeper analysis as follows.
-For the evaluation of the sensitivity of VideoLLMs towards video content, the performance on the detailed description matching of Interleave is much lower than the result on the binary classification task, which suggests that VideoLLMs may follow the question as the cue to search rare events.
-However, they are prone to overlook rare events in their detailed descriptions when no cues are provided.
-Meanwhile, VideoLLMs have difficulty in processing both binary classification and detailed description matching on Interleave.
-This indicates the over-reliance of VideoLLMs on language priors.
-Notably, VideoLLMs show worse capacity on the description matching of Interleave than Entire, which proves that common events can strongly induce VideoLLMs to generate hallucinations than with only rare events.In Misleading, despite their high accuracy on description matching, results on binary classification indicate that VideoLLMs can be affected by questions and generate event hallucinations.
-Figure 7: Accuracy of TCD on binary classification of EventHallusion with different numbers of sampled frames in the distorted video.
-Base indicates the results of standard decoding with 32 frames sampled.
-Evaluating TCD on EventHallusion
-VideoLLM Baselines.
-To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video (Li et al.
-2024a), VideoChat2 (Li et al.
-2024b), and Vila (Lin et al.
-2024) as the foundation backbones.
-LLaVA-NeXT-Video extends the ability of LLaVA (Liu et al.
-2024a) based on the projection layer design, while VideoChat2 adopts the Q-Former (Li et al.
-2023a) to bridge the gap between videos and LLM.
+As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations.
+In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories.
+A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events.
+This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model.
+(2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
+5.2 TCD Evaluation
+5.2.1 VideoLLM Baselines.
+To evaluate the effectiveness of our TCD in addressing event hallucinations, we adopt LLaVA-NeXT-Video [19], VideoChat2 [21], and VILA [26] as the foundation backbones.
+LLaVA-NeXT-Video extends the ability of LLaVA [28] based on the projection layer design, while VideoChat2 adopts the Q-Former [20] to bridge the gap between videos and LLM.
 VILA, on the other hand, employs enhanced visual-language pre-training for better video comprehension.
 Based on these MLLMs, we explore the performance of TCD across different VideoLLM architectures and designs on EventHallusion.
-Implementation Details.
+5.2.2 Implementation Details.
 We adopt different numbers of extracted frames to construct the original and distorted visual input: 32 and 8 frames for LLaVA-NeXT-Video, 12 and 8 frames for VILA, and 16 and 4 frames for VideoChat2.
-We follow (Leng et al.
-2024) and set $\alpha=1$, $\beta=0.1$.
 We use greedy search as the basic decoding paradigm.
-Main Results
-We exhibit the results of the binary classification task on EventHallusion in Tab.
-2.
-As shown in Tab.
-2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms.
-However, while TCD shows consistent and notable improvements on Interleave and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed.
+5.2.3 Main Results Analysis
+We exhibit the results of the binary classification task on EventHallusion in Tab.2.
+As shown in Tab.2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms.
+However, while TCD shows consistent and notable improvements on Mix and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed.
 This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
-Effect of Sampled Frames in Distorted Video.
+5.2.4 Ablations on Contrastive Sample Constructions.
+We also discuss some different alternative strategies when constructing the counterpart of the original video for performing our TCD method as shown in Tab.3.
+Specifically, we compare our base TCD method with adding Gaussian noises to video frames when employing TCD method (denoted as TCD+noises).
+As results have shown in Tab.3, While both TCD and TCD+noises demonstrate improvement on regular decoding.
+However, TCD+noises does not exhibit large improvements on simply using TCD method.
+Table 3: Ablation study of different strategies when constructing the contrastive counterpart of the original video. The best results of each category are bolded.
+Method Entire Mix Misleading Overall
+Baseline 49.12 70.47 71.57 64.79
+TCD 50.88 73.58 74.50 67.48
+TCD+noise 51.75 73.58 74.50 67.73
+Table 4: Ablation study of $\alpha$ in our TCD. The best results of each category are bolded.
+$\alpha$ Entire Mix Misleading Overall
+0.25 50.00 73.56 73.53 66.99
+0.5 50.88 73.56 74.51 67.48
+0.75 50.88 71.50 74.51 66.50
+1.0 49.12 64.25 74.51 62.59
+Table 5: Ablation study of $\beta$ in our TCD. The best results of each category are bolded.
+$\beta$ Entire Mix Misleading Overall
+0.1 50.88 73.58 74.51 67.48
+0.2 50.88 73.58 74.51 67.48
+0.5 50.88 74.09 74.51 67.73
+0.5 50.00 74.61 74.51 67.73
+5.2.5 Ablations on Hyper-parameters of TCD
+We conduct ablation studies to investigate the effects of hyper-parameters $\alpha$ and $\beta$ on our TCD method as shown in Tab.4 and Tab.5 For the first hyper-parameter $\alpha$, with the growth of $\alpha$, TCD applies a more fierce contrast between normal and distorted video input.
+As results are shown in Tab.4, when $\alpha$ is 0.5, TCD exhibits best performance.
+While for $\beta$, the threshold used to control the token set for TCD does not exhibit a strong impact on the performance of TCD on our EventHallusion Benchmark.
+5.2.6 Effect of Sampled Frames in Distorted Video.
 We further explore the effect of sampled frames in distorted video on our TCD.
@@
 We conduct this experiment on the binary classification task of EventHallusion.
-As results have shown in Figure 7, TCD maintains a steady improvement in the range of 4 to 16 sampled frames.
+As results have shown in Tab.6, TCD maintains a steady improvement in the range of 4 to 16 sampled frames.
 However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates.
@@
 This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
-Conclusion
+Table 6: Ablation study of extracted video frames in the distorted group. The best results of each category are bolded.
+Frames Entire Mix Misleading Overall
+1 48.26 68.39 74.51 64.30
+4 50.00 73.06 76.47 67.48
+8 50.88 73.58 74.51 67.48
+16 50.88 73.06 74.51 66.75
+Figure 6: Qualitative Results of the TCD method on our EventHallusion Benchmark.
+5.3 Qualitative Results
+We exhibit an example of our question in EventHallusion.
+As shown in Fig.6, VideoLLM can be affected and generate hallucination answers, while our TCD method can alleviate this problem.
+Due to the page limit, we provide more examples in our suuplementary materials.
+6 Conclusion
 In this paper, we systematically investigated the event-related hallucination problem of existing VideoLLMs.
-To provide a reliable evaluation benchmark to inspect such hallucination degree of VideoLLMs, we curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves.
-In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems of existing VideoLLMs in a lightweight plug-and-play manner.
-We hope our research can benefit further research in the field of reducing the hallucinations of VideoLLMs.
-Appendix A A. Automatic Evaluation For Detailed Description Matching
-We apply GPT-4o to evaluate the detailed description matching task of our Entire, Interleave, and Misleading categories respectively.
+We curated EventHallusion, which consists of multi-domain videos and annotated questions, both of which are specifically designed to trigger the priors encapsulated in the VideoLLMs themselves.
+In addition, we also proposed a simple yet effective method, called Temporal Contrastive Decoding (TCD), to alleviate the event-related hallucination problems.
+Supplementary Material
+1 Comparison with other video hallcination benmarks.
+We exhibit the comparison between HallusionBench, VideoHallucer, and our EventHallusion in Figure 7.
+As shown in Figure 7, compared to HallusionBench and Videohallucer emphasize assessing the models’ sensitivity to events’ chronological orders, our proposed EventHallusion focuses on analyzing the event-related language prior and vision-language correlation bias encapsulated in existing VideoLLMs.
+2 Automatic Evaluation For Detailed Description Matching
+We apply GPT-4o to evaluate the detailed description matching task of our Entire, Mix, and Misleading categories respectively.
 For Entire, we prompt GPT-4o to judge if VideoLLMs’ detailed description of the video is consistent with the ground-truth rare event.
-And for Interleave, we prompt GPT4o to focus on the occurrence of the interleaved rare event.
+And for Mix, we prompt GPT4o to focus on the occurrence of the rare-common mixed event.
 Finally, for Misleading, we prompt GPT4o to assess the consistency of the ground-truth event and search for any unrelated events.
@@
 The detailed prompts are shown in Figure  9, 10, 11, respectively.
+Figure 7: Comparison between HallusionBench, VideoHallucer, and our EventHallusion.
+We use Green to highlight the content of the original video/question, and red to denote refactored video/question content to assess the model’s event-related hallucinations.
 Figure 8: System prompt used in all automatic evaluation.
-Appendix B B. Qualitative Examples
-Data examples of EventHallusion.
-We present examples of Entire, Interleave, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs.
+3 Qualitative Examples
+3.1 Data examples of EventHallusion.
+We present examples of Entire, Mix, and Misleading in Figure 12, 13, 14, incorporating responses of VideoLLMs.
 As shown in these examples, VideoLLMs are prone to generate hallucinated answers under all of our settings, indicating their deficiency in addressing event hallucinations.
-Examples of GPT-4o’s Evaluation.
-We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Interleave, and Misleading respectively in Figure 15, 16, 17.
+3.2 Examples of GPT-4o’s Evaluation.
+We present examples of GPT-4o’s evaluation of detailed description matching task on Entire, Mix, and Misleading respectively in Figure 15, 16, 17.
 As shown in these examples, GPT-4o have the capacity to judge the accuracy of video events.
-Examples of TCD.
+3.3 Examples of TCD.
 We present examples of our TCD in mitigating video event hallucinations.
@@
 Figure 9: Prompt used for detailed description matching in Entire.
-Figure 10: Prompt used for detailed description matching in Interleave.
+Figure 10: Prompt used for detailed description matching in Mix.
 Figure 11: Prompt used for detailed description matching in Misleading.
 Figure 12: Examples of Entire with responses of VideoLLMs.
-Figure 13: Examples of Interleave with responses of VideoLLMs.
+Figure 13: Examples of Mix with responses of VideoLLMs.
 Figure 14: Examples of Misleading with responses of VideoLLMs.
 Figure 15: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Entire.
-Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Interleave.
+Figure 16: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Mix.
 Figure 17: Examples of the evaluation of GPT-4o on the detailed description of VideoLLMs of Misleading.